In doing a load of clothes, a clothes drier uses 16 A of current at 240 V for 42 min. A personal computer, in contrast, uses 2.9 A of current at 120 V. With the energy used by the clothes drier, how long (in hours) could you use this computer to "surf" the Internet?
Answer #1

power used by cloth drier = VI = 16*240 = 3840 W

total energy = 3840*42*60 = 9676800 J

power requirment of computer = 2.9*120 = 348 W

time for which the computer can be run = 9676800 / 348 = 27806.9 s = 7.72 hours

so, the computer can be run for 7.72 hours